OpenStudy (igigighjkl):

Jade used Fraction 3 over 4 yard of fabric to make a scarf. Can she make 2 of these scarves with Fraction 1 and 4 over 5 yards of fabric, and why? No, because the quotient of Fraction 3 over 4 ÷ Fraction 1 and 4 over 5 is 1 and 7 over 20 No, because the quotient of Fraction 1 and 4 over 5 ÷ Fraction 3 over 4 is 1 and 7 over 20 Yes, because the quotient of Fraction 3 over 4 ÷ Fraction 1 and 4 over 5 is 2 and 2 over 5 Yes, because the quotient of Fraction 1 and 4 over 5 ÷ Fraction 3 over 4 is 2 and 2 over 5 Note I Only WANT HELP!!!

OpenStudy (shiburinga):

What grade is this? :P

OpenStudy (igigighjkl):

6th

OpenStudy (shiburinga):

:O No wonder I remember it XD

OpenStudy (shiburinga):

Connections Academy or no?

OpenStudy (igigighjkl):

I'm on FLVS.

OpenStudy (welshfella):

in order to make 2 scarves she needs 2 * 3/4 yards of fabric This = 2*3 / 4 = 6/4 = 1 1/2 yards of fabric

OpenStudy (igigighjkl):

?

OpenStudy (shiburinga):

Mhm. What lesson? (My sister is in the 6th grade)

OpenStudy (welshfella):

Now the question is have we got more that 1 1/2 yards of fabric?

OpenStudy (igigighjkl):

Good. Math module 4 lesson 10.

OpenStudy (shiburinga):

Maaaaan I can't find it! D:< Stupid lessons!

OpenStudy (igigighjkl):

4.10 segment practice exam.

OpenStudy (welshfella):

- so that means the answer is YES Now you have to find which on of the last 3 choices is correct

OpenStudy (igigighjkl):

??

OpenStudy (igigighjkl):

anybody.

OpenStudy (igigighjkl):

@pooja195 please help.

OpenStudy (welshfella):

Hint divide 1 4/5 by 3/4 and see if is greater than 2

OpenStudy (shiburinga):

Sorry I can't find it. Must have changed it or something. Mhm, @welshfella is right :p

OpenStudy (igigighjkl):

0.15 is the awnser for that

OpenStudy (welshfella):

first convert 1 / 4/5 to 9 /5 then divide this by 3/4

OpenStudy (igigighjkl):

???

OpenStudy (welshfella):